1. if the area of rhombus is 48 cm^2 . and one diagonal is 6cm. find its altitude. i want answer.
 
​2. NAME A POLYHEDRON THAT HAVE 4  VERTICES 6 FACES NAD 8 EDGES
 
 

Grade:8

1 Answers

RAGHAV SHARMA
8 years ago

 

 

 1 ANS  Area of a rhombus =  (1/2) d1*d2 ....where d1, d2 are the diagonals......so  let d1  = 6   and we have

 

48 =(1/2) 6 *d2   

 

48 = 3 * d2    divide both sides by 3

 

16 = d2

 

And since the diagonals bisect each other at right angles, they form two legs of a right triangle.....so...the length of the base = √[8^2 + 3^2]  = √[64 + 9]  = √73

 

And the area can also be expressed as  a*b  where a is the altitude and b is the base....so we have

 

48 = a * b

 

48 = a * √73   

 

altitude  = 48 / √73  = about 5.618  cm

 

 

2 NOT POSSIBLE STILL IT VERIFY EUCLEAR FORMULA